OpenStudy (anonymous):

Using the equations and enthalpy values provided, which mathematical expression can be used to determine the unknown enthalpy change represented by x? 3 C + 3 O2 yields 3 CO2 deltaH = x 4 H2 + 2 O2 yields 4 H2O deltaH = - 1144 kJ C3H8 + 5 O2 yields 3 CO2 + 4 H2O deltaH = -2222 kJ C3H8 yields 3 C + 4 H2 deltaH = 104 kJ 104 = (-2222) - [x - (-1144)] 104 = (-2222) - [x + (-1144)] 104 = (-2222) + [(-1144) - x] 104 = (-2222) - [(-1144) - x] @aaronq think you could help please?

OpenStudy (anonymous):

I'll give a medal :)

OpenStudy (anonymous):

@SolomonZelman Do you think you could possibly help me? I understand if you're busy but I just dont understand this question at all.

OpenStudy (solomonzelman):

I am really bad at chemistry. I used to be very good at it, but I forgot it since 10th grade, and my maximum capability in science is answering some biology questions about Punnet squares. Sorry -:(

OpenStudy (anonymous):

@SolomonZelman It's all good. @aaronq do you know any?

OpenStudy (aaronq):

yep, i'm workin on it hold up

OpenStudy (anonymous):

ok thanks.

OpenStudy (aaronq):

This concerns Hess's Law. You have to arrange equations (1) through (3) to (4) to match the top equation (so that everything but what is there cancels out. (1) \( 3 C + 3 O_2 \rightarrow 3 CO_2 ~~~~\Delta H_1 = x\) (2) \(4 H_2 + 2 O_2 \rightarrow 4 H_2O~~~~ \Delta H_2= - 1144 kJ\) (3) \(C_3H_8 + 5 O_2 \rightarrow 3 CO_2 + 4 H_2O ~~~~\Delta H_3 = -2222 kJ\) (4) \(C_3H_8 \rightarrow 3 C + 4 H_2 ~~~\Delta H_4= 104 kJ\) So we reverse (2), and we multiply their enthalpy change by -1 (1) \( 3 C + 3 O_2 \rightarrow 3 CO_2 ~~~~\Delta H_1 = x\) (2) \( 4 H_2O\rightarrow4 H_2 + 2 O_2~~~~ \Delta H_2= 1144 kJ\) (3) \(C_3H_8 + 5 O_2 \rightarrow 3 CO_2 + 4 H_2O ~~~~\Delta H_3 = -2222 kJ\) -------------------------------------------------- (4) \(C_3H_8 \rightarrow 3 C + 4 H_2 ~~~\Delta H_4= 104 kJ\) Then we have \(\Delta H_4=\Delta H_4+\Delta H_4+\Delta H_4\)

OpenStudy (aaronq):

i made a mistake, you are to reverse (1) not (2). the rest is fine though, do you get the concept?

OpenStudy (anonymous):

Still working on it?

OpenStudy (anonymous):

Don't mean to rush you or anything

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What do you mean reverse it?

OpenStudy (anonymous):

omg I'm so sorry my connection screwed up and I didn't see your reply

OpenStudy (aaronq):

Reverse the equation, like A -> B to B -> A. When you do that you reverse the sign of \(\Delta H\)

OpenStudy (anonymous):

What is it supposed to look like once you reverse it though? I don't get it

OpenStudy (anonymous):

Also, I'm still confused over what the answer would be. Is it B?

OpenStudy (aaronq):

Sorry, this is too much to type out, and i have to step out. I think choice 4 is correct. I can help you solve it later if you want to, just tag me.

OpenStudy (anonymous):

Thanks.

OpenStudy (aaronq):

SO i double checked it and it's A. sorry about that, i can type up a solution if you're still confused.